# Provable test-time adaptivity and distributional robustness of in-context learning

### v1

@@ 3.1 Transformers
In this section, we formally define the classes of (encoder-only) Transformers that we will consider. For a vector $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, let $\operatorname{\mathsf{Softmax}}(v)\in\mathbb{R}^{d}$ denote the vector whose $j$th entry is given by $\exp(v_{j})/\sum_{\ell=1}^{d}\exp(v_{\ell})$. Often, we will apply the $\operatorname{\mathsf{Softmax}}$ function to matrices, by which we mean that the $\operatorname{\mathsf{Softmax}}$ function is applied row-wise. Given $d_{\mathrm{model}}\geq d+2$, called the model dimension, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a query $x\in\mathcal{X}$, the input matrix is defined as
$Z_{\mathrm{in}}\equiv Z_{\mathrm{in}}(D_{n},x)\coloneqq\begin{pmatrix}x_{1}^{\top}&y_{1}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}^{\top}&y_{n}&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&0\\ x^{\top}&0&\bm{0}_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$ (6)
where the last column of $Z_{\mathrm{in}}$ is our positional encoding used to identify the query, and we pad each row with zeros to make them $d_{\mathrm{model}}$-dimensional. A Transformer iteratively applies attention layers and feed forward network (FFN) layers to the input matrix $Z_{\mathrm{in}}$. Letting $N\in\mathbb{N}$ denote an upper bound on the context length, our universal approximation theory (Theorem 2) holds uniformly for all $n\in[N]$, so it is convenient to consider Transformers as functions from [^2] The space $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ is equipped with the disjoint union topology and its corresponding Borel $\sigma$-algebra. $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ to $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$. We now define attention layers.
@@
$\mathsf{FFN}_{\theta_{\mathrm{ffn}}}(Z)\coloneqq Z+\rho(ZW_{1}+\bm{1}_{n+1}v^{\top})W_{2}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}},$
@@
Definition 4 (Transformers).
We define $\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$ to be the set of all Transformers $\mathsf{TF}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ of the form
@@
For a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}$ and a query $x$, we will slightly abuse our notation by writing $\mathsf{TF}(x,D_{n})\coloneqq\mathsf{TF}\bigl(Z_{\mathrm{in}}(D_{n},x)\bigr)$ where the input matrix $Z_{\mathrm{in}}(D_{n},x)$ is defined by (6). Finally, we define $\mathsf{Read}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\mathbb{R}$ by $\mathsf{Read}(Z)\coloneqq Z_{n+1,d+1}$ for $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ and $n\in[N]$. Thus, $\mathsf{Read}\circ\mathsf{TF}(x,D_{n})$ can be interpreted as the predicted value of the response at the query $x$ based on examples $D_{n}$.
@@ 3.2 Universal approximation and learnability
$\max_{n\in[N]}\mathbb{E}_{P\sim\pi,(\mathcal{D}_{n},X,Y)|P\sim P^{\otimes(n+1)}}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\epsilon.$
@@
This proposition confirms formally the intuition provided after Proposition 1 that the first term in the decomposition (5) can be made negligible.
@@ 4.1 In-context learning for Besov functions
We remark that the Besov space $B_{\infty,\infty}^{\alpha}([0,1]^{d})$ coincides with the Hölder space $H^{\alpha}([0,1]^{d})$ for $\alpha\notin\mathbb{N}$, and it contains $H^{\alpha}([0,1]^{d})$ for $\alpha\in\mathbb{N}$; see Appendix D.1 for more details. Moreover, there exists $C_{\alpha,S,d}>0$ such that $\sup_{g\in\mathcal{G}_{\alpha}}\|g\|_{\infty}\leq C_{\alpha,S,d}$; see (54).
@@
$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+d)}$
@@
The Transformer $\widetilde{f}_{T}$ in Theorem 5 is pretrained on the mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$ with different smoothness levels, whereas the test distribution $\mu$ consists of random regression functions of a fixed smoothness level $\beta$, with a distribution shift such that $\chi^{2}(\mu,\pi_{\beta})\leq\kappa$. When we evaluate the performance of $\widetilde{f}_{T}$ under the test distribution $\mu$, Theorem 5 shows that the ICL excess $\mu$-risk of $\widetilde{f}_{T}$ is adaptive to the unknown smoothness $\beta$, in the sense that its rate of convergence $n^{-2\beta/(2\beta+d)}$ is optimal even if $\beta$ were known; see Theorem 8. Moreover, this risk bound holds uniformly over test distributions in a $\chi^{2}$-divergence ball around $\pi_{\beta}$. Thus, although the Transformer is pretrained on tasks of different levels of difficulties, it is able to achieve faster (and optimal) rates of convergence on easier tasks (without knowing the difficulty of the tasks) and is robust to distribution shift. Finally, we remark that $d_{\mathrm{model}}^{\circ},d_{\mathrm{ffn}}^{\circ}$ and $T$ should increase with $\kappa$ in order to keep $C$ independent of $\kappa$. Thus, for a Transformer with fixed architecture, fine tuning for downstream tasks may still be beneficial in cases where (i) the Transformer is not large enough, (ii) the amount of pretraining data is not sufficient, or (iii) there is significant distribution shift ($\kappa$ is large).
@@ 4.2 In-context learning for multi-index models
To avoid the curse of dimensionality in the previous section, we will now consider multi-index models with [^4] Again, our results in this section extend to any compact domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ by a scaling argument. $\mathcal{X}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}\eqqcolon\mathbb{B}^{d}$ and $\mathcal{Y}=\mathbb{R}$. For $p\in[d]$, let $V_{p}(\mathbb{R}^{d})\coloneqq\bigl\{U\in\mathbb{R}^{d\times p}:U^{\top}U=I_{p}\bigr\}$ denote the set of all projection matrices from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ (also called the Stiefel manifold). We will consider regression functions on $\mathbb{B}^{d}$ of the form $x\mapsto g\bigl((U^{\top}x+\bm{1}_{p})/2\bigr)$, where $U\in V_{p}(\mathbb{R}^{d})$ for some effective dimension $p\in[d]$ and Borel measurable $g:[0,1]^{p}\to\mathbb{R}$. Note that we translate and scale $U^{\top}x$ to ensure that it belongs to $0,1]^{p}$.
@@
$\widetilde{g}_{\alpha,p}(x)\coloneqq\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+\bm{1}_{p}}{2}\biggr),$ (11)
@@
Next, let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ with the property that there exist $c_{0}>0$ and a non-empty Euclidean ball $A\subseteq\mathbb{B}^{d}$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. We then define $\pi_{\alpha,p}$ to be the distribution of the random measure $P_{\widetilde{g}_{\alpha,p}}$, where the randomness comes from the random regression function $\widetilde{g}_{\alpha,p}\sim\widetilde{\pi}_{\alpha,p}$, and let $\mathcal{P}_{\alpha,p}\coloneqq\mathrm{supp}(\pi_{\alpha,p})$. For a finite set $\mathcal{A}^{\prime}\subseteq(0,S)\times[d]$ and a sequence of positive weights $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$ such that $\sum_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}=1$, our pretraining prior $\pi$ for multi-index models is defined to be the mixture distribution
@@
Thus $\pi$ generates random multi-index regression tasks, where the smoothness and effective dimensions of the regression functions are randomly drawn from $\mathcal{A}^{\prime}$ with probabilities given by $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$. Finally, we define $\mathcal{P}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{P}_{\alpha,p}$ and $\mathcal{G}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{G}_{\alpha,p}$. Proposition 6 is the analogue for multi-index models of Proposition 4.
@@
$\mathbb{E}_{(\mathcal{D}_{n},X,Y)\sim P_{0}^{\otimes(n+1)}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+r)}$
@@
Combining Propositions 13 and 6, we can now conclude that a pretrained Transformer is adaptive to the effective dimension of the multi-index models, as well as adapting to the smoothness of the regression function and being robust to distribution shift.
@@ 4.3 Lower bounds on the ICL excess risk
Acknowledgement: The research of the first and third author was supported by RJS’s European Research Council Advanced Grant 101019498.

### v2

@@ 3.1 Transformers
In this section, we formally define the classes of (encoder-only) Transformers that we will consider. Given $d_{\mathrm{model}}\geq d+2$, called the model dimension, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ and a query $x\in\mathcal{X}$, the input matrix $Z_{\mathrm{in}}\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ is defined as
$Z_{\mathrm{in}}\equiv Z_{\mathrm{in}}(D_{n},x)\coloneqq\begin{pmatrix}x_{1}^{\top}&y_{1}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ \vdots&\vdots&\vdots&\vdots\\ x_{n}^{\top}&y_{n}&0_{d_{\mathrm{model}}-d-2}^{\top}&0\\ x^{\top}&0&0_{d_{\mathrm{model}}-d-2}^{\top}&1\end{pmatrix}$ (5)
where the last column of $Z_{\mathrm{in}}$ is our positional encoding used to identify the query, and we pad each row with zeros to make them $d_{\mathrm{model}}$-dimensional. For a vector $v=(v_{1},\ldots,v_{d})^{\top}\in\mathbb{R}^{d}$, let $\operatorname{\mathsf{Softmax}}(v)\in\mathbb{R}^{d}$ denote the vector whose $j$th entry is given by $\exp(v_{j})/\sum_{\ell=1}^{d}\exp(v_{\ell})$. Often, we will apply the $\operatorname{\mathsf{Softmax}}$ function to matrices, by which we mean that the $\operatorname{\mathsf{Softmax}}$ function is applied row-wise. A Transformer iteratively applies attention layers and feed forward network (FFN) layers to the input matrix $Z_{\mathrm{in}}$. Letting $N\in\mathbb{N}$ denote an upper bound on the context length, our universal approximation theory (Theorem 2) holds uniformly for all $n\in[N]$, so it is convenient to consider Transformers as functions from [^2] The space $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ is equipped with the disjoint union topology and its corresponding Borel $\sigma$-algebra. $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ to $\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$. We now define attention layers.
@@
$\mathsf{FFN}_{\theta_{\mathrm{ffn}}}\!(Z)\!\coloneqq\!Z\!+\!\rho(ZW_{1}\!+\!1_{n+1}v^{\top})W_{2}\in\mathbb{R}^{(n\!+\!1)\times d_{\mathrm{model}}},$
@@
Definition 4 (Transformer).
Let $\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$ be the set of all Transformers $\mathsf{TF}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ of the form
@@
For a Transformer $\mathsf{TF}\in\mathcal{F}_{\mathrm{TF}}(L,H,d_{\mathrm{model}},d_{\mathrm{ffn}})$, examples $D_{n}=(x_{i},y_{i})_{i=1}^{n}$ and a query $x$, we will slightly abuse our notation by writing $\mathsf{TF}(x,D_{n})\coloneqq\mathsf{TF}\bigl(Z_{\mathrm{in}}(D_{n},x)\bigr)$ where the input matrix $Z_{\mathrm{in}}(D_{n},x)$ is defined by (5). Finally, we define [^3] More generally, $\mathsf{Read}(Z)$ can be an affine transformation of the last row of $Z$. $\mathsf{Read}:\bigcup_{n=1}^{N}\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}\to\mathbb{R}$ by $\mathsf{Read}(Z)\coloneqq Z_{n+1,d+1}$ for $Z\in\mathbb{R}^{(n+1)\times d_{\mathrm{model}}}$ and $n\in[N]$. Thus, $\mathsf{Read}\circ\mathsf{TF}(x,D_{n})$ can be interpreted as the predicted value of the response at the query $x$ based on examples $D_{n}$.
@@ 3.2 Universal Approximation and Learnability
$\max_{n\in[N]}\mathbb{E}_{\pi}\mathbb{E}_{P}\Bigl\{\bigl(\mathsf{Read}\circ\mathsf{TF}(X,\mathcal{D}_{n})-g_{\pi}(X,\mathcal{D}_{n})\bigr)^{2}\Bigr\}\leq\epsilon.$
@@
This proposition confirms formally the intuition provided after Proposition 1 that the first term in the decomposition (4) can be made negligible. In other words, Transformers can learn the posterior regression function from data, without knowing the pre-training prior.
@@ 4.1 In-context Learning for Besov Functions
We remark that the Besov space $B_{\infty,\infty}^{\alpha}([0,1]^{d})$ coincides with the Hölder space $H^{\alpha}([0,1]^{d})$ for $\alpha\notin\mathbb{N}$, and it contains $H^{\alpha}([0,1]^{d})$ for $\alpha\in\mathbb{N}$; see Appendix D.1 for more details. Moreover, there exists $C_{\alpha,S,d}>0$ such that $\sup_{g\in\mathcal{G}_{\alpha}}\|g\|_{\infty}\leq C_{\alpha,S,d}$; see (56). See Figure 2 for realisations of random Besov functions with $\alpha=2$ and $4$.
@@
$\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+d)}$
@@
The Transformer $\widetilde{f}_{T}$ in Theorem 5 is pretrained on the mixture distribution $\pi=\sum_{\alpha\in\mathcal{A}}\lambda_{\alpha}\pi_{\alpha}$ with different smoothness levels, whereas the test distribution $\mu$ consists of random regression functions of a fixed smoothness level $\beta$, with a distribution shift such that $\chi^{2}(\mu,\pi_{\beta})\leq\kappa$. When we evaluate the performance of $\widetilde{f}_{T}$ under the test distribution $\mu$, Theorem 5 shows that the ICL excess $\mu$-risk of $\widetilde{f}_{T}$ is adaptive to the unknown smoothness $\beta$, in the sense that its rate of convergence $n^{-2\beta/(2\beta+d)}$ is optimal even if $\beta$ were known; see Theorem 8. Moreover, this risk bound holds uniformly over test distributions in a $\chi^{2}$-divergence ball around $\pi_{\beta}$. Thus, although the Transformer is pretrained on tasks of different levels of difficulties, it is able to achieve faster (and optimal) rates of convergence on easier tasks (without knowing the difficulty of the tasks) and is robust to distribution shift. Finally, we remark that $d_{\mathrm{model}}^{\circ},d_{\mathrm{ffn}}^{\circ}$ and $T$ should increase with $\kappa$ in order to keep $C$ independent of $\kappa$. Thus, for a Transformer with fixed architecture, fine tuning for downstream tasks may still be beneficial in cases where (i) the Transformer is not large enough, (ii) the amount of pretraining data is not sufficient, (iii) there is significant distribution shift ($\kappa$ is large), or (iv) there is shift in the support of the pretraining prior and the test distribution.
@@ 4.2 In-context Learning for Multi-index Models
To avoid the curse of dimensionality in the previous section, we will now consider multi-index models with [^5] Again, our results in this section extend to any compact domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ by a scaling argument. $\mathcal{X}=\mathbb{B}^{d}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$ and $\mathcal{Y}=\mathbb{R}$. For $p\in[d]$, let $V_{p}(\mathbb{R}^{d})\coloneqq\bigl\{U\in\mathbb{R}^{d\times p}:U^{\top}U=I_{p}\bigr\}$ denote the set of all projection matrices from $\mathbb{R}^{d}$ to $\mathbb{R}^{p}$ (also called the Stiefel manifold). We will consider regression functions on $\mathbb{B}^{d}$ of the form $x\mapsto g\bigl((U^{\top}x+1_{p})/2\bigr)$, where $U\in V_{p}(\mathbb{R}^{d})$ for some effective dimension $p\in[d]$ and Borel measurable $g:[0,1]^{p}\to\mathbb{R}$. Note that we translate and scale $U^{\top}x$ to ensure that it belongs to $0,1]^{p}$.
@@
$\widetilde{g}_{\alpha,p}(x)\coloneqq\widetilde{g}_{\alpha}^{(p)}\biggl(\frac{(U^{(p)})^{\top}x+1_{p}}{2}\biggr),$ (10)
@@
Next, let $P_{X}$ be a distribution on $\mathbb{B}^{d}$ with the property that there exist $c_{0}>0$ and a non-empty Euclidean ball $A\subseteq\mathbb{B}^{d}$ such that $P_{X}(A_{0})\geq c_{0}\mathrm{Vol}_{d}(A_{0})$ for all measurable $A_{0}\subseteq A$. We then define $\pi_{\alpha,p}$ to be the distribution of the random measure $P_{\widetilde{g}_{\alpha,p}}$, where the randomness comes from the random regression function $\widetilde{g}_{\alpha,p}\sim\widetilde{\pi}_{\alpha,p}$, and let $\mathcal{P}_{\alpha,p}\coloneqq\mathrm{supp}(\pi_{\alpha,p})$. We remark that we index difficulty by the pair of parameters $(\alpha,p)$ in this section. For a finite set $\mathcal{A}^{\prime}\subseteq(0,S)\times[d]$ and a sequence of positive weights $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$ such that $\sum_{(\alpha,p)\in\mathcal{A}^{\prime}}\lambda_{\alpha,p}=1$, our pretraining prior $\pi$ for multi-index models is defined to be the mixture distribution
@@
Thus $\pi$ generates random multi-index regression tasks, where the smoothness and effective dimensions of the regression functions are randomly drawn from $\mathcal{A}^{\prime}$ with probabilities given by $(\lambda_{\alpha,p})_{(\alpha,p)\in\mathcal{A}^{\prime}}$. Finally, we define $\mathcal{P}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{P}_{\alpha,p}$ and $\mathcal{G}\coloneqq\bigcup_{(\alpha,p)\in\mathcal{A}^{\prime}}\mathcal{G}_{\alpha,p}$. Proposition 6 below is the analogue for multi-index models of Proposition 4.
@@
$\mathbb{E}_{P_{0}}\bigl[\bigl\{g_{\pi}(X,\mathcal{D}_{n})-g^{\circ}(X)\bigr\}^{2}\bigr]\leq Cn^{-2\beta/(2\beta+r)}$
@@
Combining Propositions 13 and 6, we can now conclude that a pretrained Transformer simultaneously adapts to both the effective dimension of the multi-index models and the smoothness of the regression function, and is robust to distribution shift.
@@ 4.3 Lower Bounds on the ICL Excess Risk
5 Simulations
We consider single-index models with $n=30$, $d=5$, $p=1$, $\sigma=0.01$ and four levels of Besov smoothness $\alpha\in\{2,\,2.5,\,3,\,4\}$. The generating mechanism of our pretraining data is as described in Section 4.2, with $C_{0}=2$, $(a_{k}^{(\alpha)}),(b_{\ell,\gamma}^{(\alpha)})\overset{\mathrm{iid}}{\sim}\mathrm{Unif}[-1,1]$ in (7) and $P_{X}$ being the uniform distribution on $\mathbb{B}^{d}$. For our Transformer, we use a GPT-2 medium-style model (Radford et al., 2019) with $L=24$, $H=16$, $d_{\mathrm{model}}=1024$, $d_{\mathrm{ffn}}=4096$ and GELU activation function. Further training details are provided in Appendix I.
At test time, we randomly generate new single-index models with fixed smoothness levels under two settings:
  • (i)
    No distribution shift: the random test functions are generated from the same distribution as the pretraining data of that particular smoothness,
  • (ii)
    With distribution shift: similarly as (i), except that the random variables $(a_{k}^{(\alpha)})$ in (7) are sampled from $\mathrm{Unif}[0,1]$ instead of $\mathrm{Unif}[-1,1]$.
See Figure 2 for random Besov functions with and without distribution shift (blue: no distribution shift; orange: with distribution shift).
Figure 3(a) shows the mean squared error (MSE) of our Transformer for different levels of $\alpha$; again, blue denotes ‘no distribution shift’ and orange denotes ‘with distribution shift’ at test time. We see that as the smoothness level $\alpha$ increases, our Transformer is able to adapt to the smoothness and achieve a faster rate of convergence, moreover, our Transformer is also robust to distribution shift at test time. Figure 3(b) plots the logarithm of the MSE against $2\alpha/(2\alpha+1)$, revealing that the log MSE is approximately linear in $2\alpha/(2\alpha+1)$. This is in line with the conclusion of Theorem 7.
(a) $\alpha=2$.
(b) $\alpha=4$.
Figure 2: Random Besov functions with smoothness $\alpha\in\{2,4\}$. Blue represents functions with no distribution shift and orange represents functions with distribution shift.
(a) MSE against $\alpha$.
(b) Log MSE against $\frac{2\alpha}{2\alpha+1}$.
Figure 3: MSE and log MSE at test time based on 100,000 random test functions. Error bars represent 90% confidence intervals.
6 Conclusion
In this work, we show that pretrained Transformers adapt to different task difficulties and are robust to distribution shifts at test time. This is proved by first showing that Transformers are universal approximators of posterior regression functions, and then establishing rates of convergence of the posterior regression functions on each task. Our results connect ICL with posterior contraction theory, and provide a rigorous explanation for the empirical success of ICL.
A natural extension of this work would be to combine our theory with the quantitative approximation theory derived in independent, concurrent work by Wakayama and Suzuki (2025) to provide more explicit values of $d_{\mathrm{model}}^{\circ}$, $d_{\mathrm{ffn}}^{\circ}$ and $T$ that guarantee the conclusions of Theorems 5 and 7. More broadly, our work opens the door to providing adaptive and distributionally robust ICL guarantees for Transformers in other settings, including heavy-tailed noise, classification and dependent data.
Acknowledgement
The research of the first and third author was supported by RJS’s European Research Council Advanced Grant 101019498.